OpenStudy (anonymous):

If the common difference in an arithmetic sequence is –3, and the first term is 7, what is the 14th term?

OpenStudy (texaschic101):

an = a1 + (n -1)d n = 14 a1 = 7 d = -3 a14 = 7 + (14 - 1)-3 a14 = 7 + (13)(-3) a14 = 7 + (-39) a14 = -32 14th term is - 32 (I think) http://www.regentsprep.org/Regents/math/algtrig/ATP2/ArithSeq.htm